# Bayesian Optimisation with Unknown Hyperparameters: Regret Bounds Logarithmically Closer to Optimal

### v1

@@ 1 Introduction
Bayesian Optimisation (BO) [15] has proven to be an efficient solution for black-box optimisation problems, finding applications across science, engineering and machine learning [12, 17, 22]. As a model-based optimisation technique, BO constructs a surrogate model of the black box function, which is typically a Gaussian Process (GP) [35]. However, to construct this surrogate, we need to specify our expectations about the smoothness of the black-box function. In the case of GP, the choice of smoothness is reflected in the selection of an appropriate length scale value for the kernel function. Selecting smaller length scales allows us to model less smooth functions and, as such, expands the class of all possible black-box functions the optimiser will consider. At the same time, it makes the convergence of the algorithm slower, due to an increase in a number of possible ‘candidate’ functions, the algorithm has to explore the space much more. As such, we wish to consider the smallest possible class of functions that still contains the black-box function we wish to solve. This translates to selecting some optimal length scale value, which is neither too short nor too long.
Figure 1: An objective function (proposed by [6]) that illustrates the importance of length scales to BO. The blue line shows a GP fit with shaded regions representing one standard deviation. The length scale value was set to the optimal value on the left and was selected by MLE on the right, based on five points represented by dots. While the optimiser with MLE length scale persistently selects suboptimal value of $x=1$, the optimiser with the optimal length scale can spot the hidden peak leading to finding the maximum at $x^{*}=0.3$.
@@
A-GP-UCB is suboptimal because it never returns to previously trialled, longer length scales. Observe that if trying a shorter length scale value does not improve function discovery, opting for a longer scale is a safer choice, preventing excessive exploration. To build an algorithm following this intuition, we could have a number of base optimisers, each utilising a different length scale value, and aggregate them into a single ‘master’ optimiser. By knowing how explorative each of the base optimisers is, the ‘master’ optimiser could select the most suitable one at each iteration, so as to balance exploration and exploitation. Within the literature of multi-armed bandit problems, a number of rules for aggregating base algorithms has been proposed [1, 3, 29], however, the performance of those ‘master’ algorithms worsens with the number of base algorithms. This prohibits us from directly applying ‘master’ algorithms to the unknown hyperparameter problem, as the length scale is a continuous parameter and has infinitely many possible values.
@@
Within this work, we extend one such aggregation scheme, called regret-balancing, to handle infinitely many learners, so that it could tackle the problem of BO with unknown hyperparameters. We propose an algorithm called Length scale Balancing GP-UCB (LB-GP-UCB), which aggregates a number of base optimisers with different length scale values and gradually introduces new base optimisers, equipped with smaller length scales. Instead of permanently decreasing the length scale value, as done by A-GP-UCB, LB-GP-UCB occasionally introduces new base learners with smaller length scale values, while still maintaining base learners with longer ones. As such, if one of the longer length scales is optimal, we will be able to recover performance close to the one of the oracle optimiser utilising that optimal length scale. Denoting the factor by which the regret bound of A-GP-UCB was away from oracle as $g(T)$, we show that LB-GP-UCB is only $\log g(T)$ away from oracle regret. We also conduct empirical evaluation and show LB-GP-UCB obtains improved regret results on a mix of synthetic and real-world benchmarks both compared to A-GP-UCB, MLE and MCMC. We show the histogram of length scale values selected by each method on one of the benchmark problems in Figure 2. $\hat{\theta}^{\star}$ represents an estimate of optimal length scale value on this problem (see §5 for details). We can see that the length scale values selected by LB-GP-UCB are close to the estimated optimal value, whereas MLE and A-GP-UCB miss this value by respectively over and under-estimating, matching our predictions. We summarise our contributions below.
@@
  • We extend our algorithm to also handle the case of unknown output scale (function norm) along the length scale
@@ 2 Problem Statement and Preliminaries
We consider the problem of maximising an unknown black-box function $f:\mathcal{X}\to\mathbb{R}$ for some compact set $\mathcal{X}\subset\mathbb{R}^{d}$. At each time step $t$, we are allowed to query a function at a selected point $\bm{x}_{t}\in\mathcal{X}$ and observe noisy feedback $y_{t}=f(\bm{x}_{t})+\epsilon_{t}$, where $\epsilon_{t}\sim\textrm{SubGauss}(\sigma_{N}^{2})$. We wish to find the optimum $\bm{x}^{\star}=\max_{\bm{x}\in\mathcal{X}}f(\bm{x})$. We define the instantaneous regret to be $r_{t}=f(\bm{x}^{\star})-f(\bm{x}_{t})$ and cumulative regret as $R_{T}=\sum_{t=1}^{T}r_{t}$, and we wish to minimise it. We assume we are given some kernel function $k^{\theta}(\bm{x},\bm{x}^{\prime})=k(\frac{\bm{x}}{\theta},\frac{\bm{x}^{\prime}}{\theta})$ parametrised by a length scale value $\theta\in\mathbb{R}^{+}$ and we denote its associated Reproducing Kernel Hilbert Space (RKHS) as $\mathcal{H}(k^{\theta})$. We assume that at least for certain values of $\theta$, the black-box function $f$ belongs to this RKHS, i.e. $\exists_{\theta\in\mathbb{R}^{+}}f\in\mathcal{H}(k^{\theta})$. Concretely, we will consider two popular types of kernels: RBF and $\nu$-Matérn, defined below for completeness:
@@
Let $f\in\mathcal{H}\left(k^{\theta}\right)$, such that $\lVert f\rVert_{k^{\theta}}\leq B$ and set $\beta_{t}^{\theta,B}=B+\sigma_{N}\sqrt{2(\gamma(k^{\theta})_{t-1}+1+\ln(1/\delta_{A}))}$, where $\gamma_{T}(k^{\theta})$ is an upper bound $\frac{1}{2}\log\lvert I+\sigma_{N}^{2}\bm{K}^{\theta^{*}}_{T}\rvert\leq\gamma_{T}(k^{\theta})$, which depends on the kernel and length scale choice. Then, with probability at least $1-\delta_{A}$, for all $\bm{x}\in\mathcal{X}$ and $t=1,\dots,T$:
@@
Let us run a GP-UCB utilising a GP with a kernel $k^{\theta}$ and the exploration bonus of $\beta_{t}^{\theta,B}=B+\sigma_{N}\sqrt{2(\gamma_{t-1}(k^{\theta})+1+\ln(1/\delta_{A}))}$ on a black-box function $f\in\mathcal{H}(k^{\theta})$ such that $\lVert f\rVert_{k^{\theta}}\leq B$. Then, with probability at least $1-\delta_{A}$, it admits the following bound on its cumulative regret $R_{T}\leq R^{\theta,B}(T)=\mathcal{O}\left(\sqrt{T}\left(B\sqrt{\gamma_{T}(k^{\theta})}+\gamma_{T}(k^{\theta})\right)\right)$.
Note that the Theorem 2.2 relies the quantity $\gamma_{T}(k^{\theta})$, also called maximum information gain (MIG). The next proposition provides bounds on $\gamma_{T}(k^{\theta})$ for RBF and $\nu$-Matérn kernel and shows the explicit dependence on length scale hyperparameter $\theta$.
Proposition 2.3 (Proposition 2 in [6]).
We have that:
  • For an RBF kernel $\gamma_{T}(k^{\theta})=\mathcal{O}\left(\frac{1}{\theta^{d}}\log(T)^{d+1}\right)$
  • For $\nu$-Matérn kernel $\gamma_{T}(k^{\theta})=\mathcal{O}\left(\frac{1}{\theta^{2\nu+d}}T^{\frac{d(d+1)}{2\nu+d(d+1)}}\log(T)\right)$
@@
The proposition above shows that the regret bound grows as $\theta$ decreases. In order for GP-UCB to enjoy regret bound from Theorem 2.2, we need to know an upper bound on the RKHS norm $B$ of the black-box function for the given kernel $k^{\theta}$, so that we can set the exploration bonus $\beta_{t}^{\theta,B}$. Inspecting the regret bound, we see that selecting the smallest $B$ (i.e. the tightest bound) is the optimal choice. Note that the same function $f(\cdot)$, can have different RKHS norms under kernels with different length scale values. As such, to obtain the optimal scaling of the regret bound, one needs to jointly optimise for $\theta$ and $B$. The optimal hyperparameters are thus $\theta^{\star},B^{\star}=\arg\min_{\theta,B\in\mathbb{R}^{+}}R^{\theta,B}(T)$ such that $\lVert f\rVert_{k^{\theta^{\star}}}\leq B^{\star}$. We assume we are given some initial $\theta_{0}\geq\theta^{\star}$ and $B_{0}\leq B^{\star}$. As explained in the introduction, in practice, those initial values could be found by maximising the marginal likelihood for a small number of initial data points. We now notice one interesting property. In case of RBF and $\nu$-Matérn kernels, let us assume, we increase the norm by $\alpha>1$ and decrease the length scale by $\beta>1$. We then get that the regret bound with those new parameters scales as follows:
$R^{\theta/\beta,\alpha B}(T)=\mathcal{O}\left(\sqrt{T}\left(\alpha\left(\frac{\beta}{\theta}\right)^{\hat{d}/2}B\sqrt{\gamma_{T}(k^{\theta})}+\left(\frac{\beta}{\theta}\right)^{\hat{d}}\gamma_{T}(k^{\theta})\right)\right),$
@@
where $\hat{d}=d$ in case of an RBF kernel and $\hat{d}=d+2\nu$ in case of a $\nu$-Matérn kernel. Since $\gamma_{T}(k^{\theta})$ is increasing in $T$, for large enough $T$ we have $B<\sqrt{\gamma_{T}(k^{\theta})}$ and any $\alpha<\left(\frac{\beta}{\theta}\right)^{d/2}$ does not affect the bound scaling. As such, whenever we decrease lengthscale by $\beta$, we can increase bound by $\left(\frac{\beta}{\theta}\right)^{d/2}$ essentially "for free". As such, we are going to use $\theta$-dependent norms in form of $B(\theta,N)=(\frac{\theta_{0}}{\theta})^{d}N$, where $N$ the norm bound under $\theta_{0}$, becomes the new hyperparameter we wish to select, instead of $B$. The optimal values of hyperparameters under this new parameterization are thus $\theta^{\star},N^{\star}=\arg\min R^{\theta,B(\theta,N)}(T)$ subject to $\lVert f\rVert_{k^{\theta^{\star}}}\leq B(\theta^{\star},N^{\star})$. Notice that $B_{0}\leq B^{\star}\leq B(\theta^{\star},N^{\star})\leq B(\theta_{0},N^{\star})=N^{\star}$ and as such, the initial value $B_{0}$ is also a lower bound on $N^{\star}$. Using this new parameterization has one important benefit, as stated by next Lemma.
Lemma 2.1 (Lemma 4 in [9]).
@@
For any $\theta<\theta^{\star}$ and $N>N^{\star}$, we have that $\lVert f\rVert_{k^{\theta}}\leq B(\theta,N)$.
@@
We will thus refer to any pair $(\theta,N)$, such that $\theta\leq\theta^{\star}$ and $N\geq N^{\star}$, as well-specified hyperparameters, as the GP-UCB admits a provable regret bound when they are used (albeit that bound might not be optimal). For simplicity, we are now going to assume that $N^{\star}$ is known and proceed with solving the problem of only one unknown hyperparameter $\theta^{\star}$. We will thus be writing $\beta^{\theta}=\beta^{\theta,B(\theta,N^{\star})}$, $\textrm{UCB}_{t}^{\theta}=\textrm{UCB}_{t}^{\theta,B(\theta,N^{\star})}$ and $R^{\theta}(T)=R^{\theta,B(\theta,N^{\star})}(T)$. However, we would like to emphasise that the algorithm we will propose throughout this paper can be extended to the case when $B^{\star}$ is also an unknown hyperparameter, which we do in Appendix D.
@@ 3 Length scale Balancing
Aggregation schemes describe a set of rules that a master algorithm should follow while coordinating a number of base algorithms. One such scheme is regret-balancing with elimination [29]. This scheme assumes each of the base algorithms comes with a suspected regret bound, that is a high-probability bound on its regret that holds if the algorithm is well-specified for the given problem, but might not hold if the learner is misspecified. The scheme always selects the base algorithm that currently has the smallest cumulative regret according to its suspected bound. It also removes base algorithms that underperform, compared to others, by more than their suspected bound, as this means their bound does not hold and, with a high probability, are misspecified. Our idea is to have each base algorithm be a GP-UCB algorithm with a different value of the length scale hyperparameter. Let us now discuss how to identify candidates for the length scale values. We propose the usage of a candidate-suggesting function $q(\cdot):\mathbb{N}\to\mathbb{R}^{+}$, such that the $i$th candidate length scale value to consider is given by $q(i)$.
@@
$q(i)=\theta_{0}e^{-i/\hat{d}},$
where $\hat{d}=d$ in case of an RBF kernel and $\hat{d}=d+2\nu$ in case of $\nu$-Matérn kernel.
We want to ensure that one of the candidates we will eventually introduce will be close to $\theta^{*}$. Let us denote $\hat{\theta}=\arg\max_{i\in\mathbb{N}\textrm{ };\,q(i)\leq\theta^{\star}}q(i)$ to be the largest length scale suggested by our candidate-suggesting function that is still smaller than $\theta^{\star}$. Observe that $f\in\mathcal{H}(k^{\hat{\theta}})$ and thus $\hat{\theta}$ is the largest well-specified length scale value among suggested candidates. As such, the regret bound of the best base learner is $R^{\hat{\theta}}(T)$ and we hope that the regret bound of a master algorithm aggregating this learner with others will be close to $R^{\hat{\theta}}(T)$. Comparing with the regret bound of the GP-UCB algorithm utilising the true optimal length scale value $R^{\theta^{\star}}(T)$, we get the result stated by the following Lemma 3.1, proven in Appendix B
@@
In case of both RBF and $\nu$-Matérn kernel we have that:
@@
This Lemma shows that the regret bound of the best of our base algorithms is only a constant factor away from the bound of the algorithm using the optimal length scale value. However, as for any $i\in\mathbb{N}$, we have $q(i)>0$, the candidate-suggesting function $q(\cdot)$ introduces infinitely many candidates. As we can only aggregate a finite number of base algorithms, we thus propose to gradually introduce new optimisers equipped with new candidate length scale values. Observe that if we stopped our quantisation at some lower bound $\theta_{L}$, then we would create a maximum of $q^{-1}(\theta_{L})$ candidates, that is, $\hat{d}\ln(\frac{\theta_{0}}{\theta_{L}})$. However, this would require us to know a sure lower bound on the optimal length scale value. Since we do not have this knowledge, we could employ a mechanism similar to A-GP-UCB, where we progressively decrease the suspected lower bound value $\theta_{L}(t)=\frac{\theta_{0}}{g(t)}$, based on some growth function $g(t)$. Observe that since $\ln\left(\frac{\theta_{0}}{\theta_{L}(t)}\right)=\ln(g(t))$, the number of candidate values grows only logarithmically with the growth function $g(t)$. Same as for A-GP-UCB, this growth function needs to be specified by the user, and we describe how this choice can be made in §5. However, we would like to emphasise that, unlike A-GP-UCB which simply sets its length scale value to $\theta_{L}(t)$, we instead introduce new learners with shorter length scale values, while still keeping the old learners with longer values. This strategy is thus more robust to the choice of the growth function, which is reflected in better scaling of the regret bound we derive later. In Algorithm 1, we present LB-GP-UCB, an algorithm employing this mechanism. We now briefly explain the logic behind its operations.
@@
11:        {# Eliminate underperforming length scale}
12:        $\Theta_{t+1}=\left\{\theta\in\Theta_{t}:L_{t}(\theta)+\frac{1}{|S_{t}^{\theta}|}\sum_{\tau\in S_{t}^{\theta}}\beta_{\tau}^{\theta}\sigma^{\theta}_{\tau-1}(\bm{x}_{\tau})\geq\max_{\theta^{\prime}\in\Theta_{t}}L_{t}(\theta^{\prime})\right\}$

### v3

@@ 1 Introduction
Bayesian Optimisation (BO) [16] has proven to be an efficient solution for black-box optimisation problems, finding applications across science, engineering and machine learning [12, 18, 23]. As a model-based optimisation technique, BO constructs a surrogate model of the black box function, which is typically a Gaussian Process (GP) [40]. However, to construct this surrogate, we need to specify our expectations about the smoothness of the black-box function. In the case of GP, the choice of smoothness is reflected in the selection of an appropriate length scale value for the kernel function. Selecting smaller length scales allows us to model less smooth functions and, as such, expands the class of all possible black-box functions the optimiser will consider. At the same time, it makes the convergence of the algorithm slower, due to an increase in the number of possible ‘candidate’ functions, the algorithm has to explore the space much more. As such, we wish to consider the smallest possible class of functions that still contains the black-box function we wish to solve. This translates to selecting some optimal length scale value, which is neither too short nor too long.
Figure 1: An objective function (proposed by [6]) that illustrates the importance of length scales to BO. The blue line shows a GP fit with shaded regions representing one standard deviation. The length scale value was set to the optimal value on the left and was selected by MLE on the right, based on five points represented by dots. While the optimiser with the MLE of length scale persistently selects a suboptimal value of $x=1$, the optimiser with the optimal length scale can spot the hidden peak leading to finding the maximum at $x^{*}=0.3$.
@@
A-GP-UCB is suboptimal because it never returns to previously trialled, longer length scales. Observe that if trying a shorter length scale value does not improve function discovery, opting for a longer scale is a safer choice, preventing excessive exploration. To build an algorithm following this intuition, we could have a number of base optimisers, each utilising a different length scale value, and aggregate them into a single ‘master’ optimiser. By knowing how explorative each of the base optimisers is, the ‘master’ optimiser could select the most suitable one at each iteration, so as to balance exploration and exploitation. Within the literature of multi-armed bandit problems, a number of rules for aggregating base algorithms have been proposed [1, 3, 33], however, the performance of those ‘master’ algorithms worsens with the number of base algorithms. This prohibits us from directly applying ‘master’ algorithms to the unknown hyperparameter problem, as the length scale is a continuous parameter and has infinitely many possible values.
@@
Within this work, we extend one such aggregation scheme, called regret-balancing, to handle infinitely many learners, so that it could tackle the problem of BO with unknown hyperparameters. We propose an algorithm called Length scale Balancing GP-UCB (LB-GP-UCB), which aggregates a number of base optimisers with different length scale values and gradually introduces new base optimisers, equipped with smaller length scales. Instead of permanently decreasing the length scale value, as done by A-GP-UCB, LB-GP-UCB occasionally introduces new base learners with smaller length scale values, while still maintaining base learners with longer ones. As such, if one of the longer length scales is optimal, we will be able to recover performance close to the one of the oracle optimiser utilising that optimal length scale. Denoting the factor by which the regret bound of A-GP-UCB was away from oracle as $g(T)$, we show that LB-GP-UCB is only $\log g(T)$ away from oracle regret. We also conduct empirical evaluation and show LB-GP-UCB obtains improved regret results on a mix of synthetic and real-world benchmarks compared to A-GP-UCB, MLE and MCMC. We show the histogram of length scale values selected by each method on one of the benchmark problems in Figure 2. $\hat{\theta}^{\star}$ represents an estimate of optimal length scale value on this problem (see §5 for details). We can see that the length scale values selected by LB-GP-UCB are close to the estimated optimal value, whereas MLE and A-GP-UCB miss this value by respectively over and under-estimating, matching our predictions. We summarise our contributions below.
@@
  • We extend our algorithm to also handle the case of unknown output scale (function norm) alongside the length scale
@@ 2 Problem Statement and Preliminaries
We consider the problem of maximising an unknown black-box function $f:\mathcal{X}\to\mathbb{R}$ on some compact set $\mathcal{X}\subset\mathbb{R}^{d}$. At each time step $t$, we are allowed to query a function at a selected point $\bm{x}_{t}\in\mathcal{X}$ and observe noisy feedback $y_{t}=f(\bm{x}_{t})+\epsilon_{t}$, where $\epsilon_{t}\sim\textrm{SubGauss}(\sigma_{N}^{2})$. We wish to find the optimum $\bm{x}^{\star}=\max_{\bm{x}\in\mathcal{X}}f(\bm{x})$. We define the instantaneous regret to be $r_{t}=f(\bm{x}^{\star})-f(\bm{x}_{t})$ and cumulative regret as $R_{T}=\sum_{t=1}^{T}r_{t}$, and we wish to minimise it. We assume we are given some kernel function $k^{\theta}(\bm{x},\bm{x}^{\prime})=k(\frac{\bm{x}}{\theta},\frac{\bm{x}^{\prime}}{\theta})$ parametrised by a length scale value $\theta\in\mathbb{R}^{+}$ and we denote its associated Reproducing Kernel Hilbert Space (RKHS) as $\mathcal{H}(k^{\theta})$. We assume that at least for certain values of $\theta$, the black-box function $f$ belongs to this RKHS, i.e. $\exists_{\theta\in\mathbb{R}^{+}}f\in\mathcal{H}(k^{\theta})$. Concretely, we will consider two popular types of kernels: RBF and $\nu$-Matérn, defined below for completeness:
@@
Let $f\in\mathcal{H}\left(k^{\theta}\right)$, such that $\lVert f\rVert_{k^{\theta}}\leq B$ and set $\beta_{t}^{\theta,B}=B+\sigma_{N}\sqrt{2(\mathcal{I}_{t-1}(k^{\theta})+1+\ln(1/\delta_{A}))}$, where $\mathcal{I}_{T}(k^{\theta})$ is an upper bound $\frac{1}{2}\log\lvert I+\sigma_{N}^{-2}\bm{K}^{\theta^{\star}}_{T}\rvert\leq\mathcal{I}_{T}(k^{\theta})$, which depends on the kernel and length scale choice. Then, with probability at least $1-\delta_{A}$, for all $\bm{x}\in\mathcal{X}$ and $t=1,\dots,T$:
@@
Note that the Theorem 2.1 relies on the quantity $\mathcal{I}_{T}(k^{\theta})$, also called maximum information gain (MIG). The next proposition, proven in Appendix A, provides bounds on $\mathcal{I}_{T}(k^{\theta})$ for RBF and $\nu$-Matérn kernel and shows the explicit dependence on length scale hyperparameter $\theta$.
Proposition 2.2.
We have that $\mathcal{I}_{T}(k^{\theta})\leq\mathcal{O}\left(\gamma_{T}(k^{\theta})\right)$:
  • For an RBF kernel $\gamma_{T}(k^{\theta})=\frac{1}{\theta^{d}}\log(T)^{d+1}$
  • For $\nu$-Matérn kernel $\gamma_{T}(k^{\theta})=\frac{1}{\theta^{d}}T^{\frac{d(d+1)}{2\nu+d(d+1)}}\log(T)^{\frac{2\nu}{2\nu+d}}$
@@
Let us run a GP-UCB utilising a GP with a kernel $k^{\theta}$ and the exploration bonus of $\beta_{t}^{\theta,B}=B+\sigma_{N}\sqrt{2(\mathcal{I}_{T}(k^{\theta}))+1+\ln(1/\delta_{A}))}$ on a black-box function $f\in\mathcal{H}(k^{\theta})$ such that $\lVert f\rVert_{k^{\theta}}\leq B$. Then, with probability at least $1-\delta_{A}$, it admits the following bound on its cumulative regret $R_{T}\leq\mathcal{O}\left(R^{\theta,B}(T)\right)$, where $R^{\theta,B}(T)=\sqrt{T}\left(B\sqrt{\gamma_{T}(k^{\theta})}+\gamma_{T}(k^{\theta})\right)$.
In our notation, note the distinction between the regret of an algorithm $R_{T}$ and the scaling of its bound $R^{\theta,B}(T)$. As the maximum regret we can possibly suffer at any time step, while optimising a function with property $\lVert f\rVert_{k^{\theta}}\leq B$, is bounded as $r_{t}\leq 2B$  [^1] This is because $f(\bm{x}^{\star})-f(\bm{x}_{t})\leq 2\lVert f\rVert_{\infty}\leq 2\lVert f\rVert_{k^{\theta}}\leq 2B$., we are going to assume the bound obeys the property $R^{\theta,B}(t+1)-R^{\theta,B}(t)\leq 2B$ for all $t=1,\dots,T-1$, as otherwise the bound can be trivially improved.
@@
In order for the bound of Theorem 2.1 to hold, we need to know the length scale $\theta$ and an upper bound on the RKHS norm $B$ of the black-box function for the given kernel $k^{\theta}$. Inspecting the regret bound together with Proposition 2.2, we see that selecting the smallest $B$ (i.e. the tightest bound) and the longest length scale $\theta$ results in the smallest $R^{\theta,B}(T)$. Note that the same function $f(\cdot)$, can have different RKHS norms under kernels with different length scale values. As such, to obtain the optimal scaling of the regret bound, one needs to jointly optimise for $\theta$ and $B$. The optimal hyperparameters are thus $\theta^{\star},B^{\star}=\arg\min_{\theta,B\in\mathbb{R}^{+}}R^{\theta,B}(T)$ such that $\lVert f\rVert_{k^{\theta^{\star}}}\leq B^{\star}$. We assume we are given some initial $\theta_{0}\geq\theta^{\star}$ and $B_{0}\leq B^{\star}$. As explained in the introduction, in practice, those initial values could be found by maximising the marginal likelihood for a small number of initial data points. We now notice one interesting property. In the case of RBF and $\nu$-Matérn kernels, if we change the length scale value from $\theta_{0}$ to $\theta$ and the norm bound from $B_{0}$ to $B$, we get that the regret bound with those new hyperparameters scales as follows:
$R^{\theta,B}(T)=\sqrt{T}\left(\left(\frac{B}{B_{0}}\right)\left(\frac{\theta_{0}}{\theta}\right)^{d/2}B_{0}\sqrt{\gamma_{T}(k^{\theta_{0}})}+\left(\frac{\theta_{0}}{\theta}\right)^{d}\gamma_{T}(k^{\theta_{0}})\right).$
@@
Since $\gamma_{T}(k^{\theta})$ is increasing in $T$, for large enough $T$ we have $B<\sqrt{\gamma_{T}(k^{\theta})}$ and any $B<\left(\frac{\theta_{0}}{\theta}\right)^{d/2}B_{0}$ does not affect the bound’s order dependence. As such, whenever we decrease lengthscale to $\theta$, we can increase norm bound by $\left(\frac{\theta_{0}}{\theta}\right)^{d/2}$ essentially "for free". As such, we are going to use $\theta$-dependent norms in the form of $B(\theta,N)=(\frac{\theta_{0}}{\theta})^{d/2}N$, where $N$ is the norm bound under $\theta_{0}$ and becomes the new hyperparameter we wish to select, instead of $B$. The optimal values of hyperparameters under this new parameterization are thus $\theta^{\star},N^{\star}=\arg\min R^{\theta,B(\theta,N)}(T)$ subject to $\lVert f\rVert_{k^{\theta^{\star}}}\leq B(\theta^{\star},N^{\star})$. Notice that $\min_{\theta\in(0,\theta_{0}]}B(\theta,N^{\star})=N^{\star}$ and as such it does not make sense to try values of $N$ smaller than $B_{0}$. Using this new parameterization brings an important benefit, as stated next.
Lemma 2.1 (Consequence of Lemma 4 in [9]).
@@
In case of RBF and Matérn kernels, for any $\theta<\theta^{\star}$ and $N>N^{\star}$, we have that $\lVert f\rVert_{k^{\theta}}\leq B(\theta,N)$.
@@
We will thus refer to any pair $(\theta,N)$, such that $\theta\leq\theta^{\star}$ and $N\geq N^{\star}$, as well-specified hyperparameters, as the GP-UCB admits a provable regret bound when they are used (albeit that bound might not be optimal). For simplicity, we are now going to assume that $N^{\star}$ is known and proceed with solving the problem of only one unknown hyperparameter $\theta^{\star}$. We will thus be writing $\beta^{\theta}_{t}=\beta^{\theta,B(\theta,N^{\star})}_{t}$, $\textrm{UCB}_{t}^{\theta}(\cdot)=\textrm{UCB}_{t}^{\theta,B(\theta,N^{\star})}(\cdot)$ and $R^{\theta}(\cdot)=R^{\theta,B(\theta,N^{\star})}(\cdot)$. However, we would like to emphasise that the algorithm we will propose throughout this paper can be extended to the case when $B^{\star}$ is also an unknown hyperparameter, which we do in Appendix F.
@@ 3 Length scale Balancing
Aggregation schemes describe a set of rules that a master algorithm should follow while coordinating a number of base algorithms. One such scheme is regret-balancing with elimination [33]. This scheme assumes each of the base algorithms comes with a suspected regret bound, which is a high-probability bound on its regret that holds if the algorithm is well-specified for the given problem, but might not hold if the learner is misspecified. The scheme always selects the base algorithm that currently has the smallest cumulative regret according to its suspected bound. This ensures that the regret of the master algorithm will not be too far from the regret of the best well-specified candidate. It also removes base algorithms that underperform, compared to others, by more than their suspected bound, as this means their bounds do not hold and, with a high probability, are misspecified.
Our idea is to use regret balancing with elimination while having each base algorithm be a GP-UCB algorithm with a different value of the length scale hyperparameter. Let us now discuss how to identify candidates for the length scale values. We propose the usage of a candidate-suggesting function $q(\cdot):\mathbb{N}\to\mathbb{R}^{+}$, such that the $i$th candidate length scale value to consider is given by $q(i)$.
@@
$q(i)=\theta_{0}e^{-i/d}.$
We want to ensure that one of the candidates we will eventually introduce will be close to $\theta^{\star}$. Let us denote $\hat{\theta}=\arg\max_{i\in\mathbb{N}\textrm{ };\,q(i)\leq\theta^{\star}}q(i)$ to be the largest length scale suggested by our candidate-suggesting function that is still smaller than $\theta^{\star}$. Observe that $\lVert f\rVert_{k^{\hat{\theta}}}\leq B(\hat{\theta},N^{\star})$ and thus $\hat{\theta}$ is the largest well-specified length scale value among suggested candidates. As such, the regret bound of the best base learner is $R^{\hat{\theta}}(T)$ and we hope that the regret bound of a master algorithm aggregating this learner with others will be close to $R^{\hat{\theta}}(T)$. Comparing with the regret bound of the GP-UCB algorithm utilising the true optimal length scale value $R^{\theta^{\star}}(T)$, we get the result stated by the following Lemma 3.1, proven in Appendix C
@@
In the case of both RBF and $\nu$-Matérn kernel, we have that:
@@
This Lemma shows that the regret bound of the best of our base algorithms is only a constant factor away from the bound of the algorithm using the optimal length scale value. However, as for any $i\in\mathbb{N}$, we have $q(i)>0$, and the candidate-suggesting function $q(\cdot)$ introduces infinitely many candidates. As we can only aggregate a finite number of base algorithms, we thus propose to gradually introduce new optimisers equipped with new candidate length scale values. Observe that if we stopped our quantisation at some lower bound $\theta_{L}$, then we would create a maximum of $q^{-1}(\theta_{L})$ candidates, that is, $d\ln(\frac{\theta_{0}}{\theta_{L}})$. However, this would require us to know a sure lower bound on the optimal length scale value. Since we do not have this knowledge, we could employ a mechanism similar to A-GP-UCB, where we progressively decrease the suspected lower bound value $\theta_{L}(t)=\frac{\theta_{0}}{g(t)}$, based on some growth function $g(t)$. Observe that since $\ln\left(\frac{\theta_{0}}{\theta_{L}(t)}\right)=\ln(g(t))$, the number of candidate values grows only logarithmically with the growth function $g(t)$. Same as for A-GP-UCB, this growth function needs to be specified by the user, and we describe how this choice can be made in §5. However, we would like to emphasise that, unlike A-GP-UCB which simply sets its length scale value to $\theta_{L}(t)$, we instead introduce new learners with shorter length scale values, while still keeping the old learners with longer values. This strategy is thus more robust to the choice of the growth function, which is reflected in better scaling of the regret bound we derive later. In Algorithm 1, we present LB-GP-UCB, an algorithm employing this mechanism. We now briefly explain the logic behind its operations.
@@
11:        {# Eliminate underperforming length scale values}
12:        $\Theta_{t+1}=\left\{\theta\in\Theta_{t}:L_{t}(\theta)+\frac{2}{|S_{t}^{\theta}|}\sum_{\tau\in S_{t}^{\theta}}\beta_{\tau}^{\theta}\sigma^{\theta}_{\tau-1}(\bm{x}_{\tau})\geq\max_{\theta^{\prime}\in\Theta_{t}}L_{t}(\theta^{\prime})\right\}$